# MEAL: A Benchmark for Continual Multi-Agent Reinforcement Learning

### v1

@@ A.2 Procedural Kitchen Generator
Given a random seed and user-selectable parameters (number of agents $n_{a}$, layout height range $[h_{\min},h_{\max}]$, layout width range $[w_{\min},w_{\max}]$, and wall-density $\rho$), the goal is to emit a solvable grid string $G$ representing the Overcooked environment.
@@ A.2.1 Notation
[…]
@@ A.2.2 Algorithm
The generator performs the following loop until a valid grid is produced (Listing A.2.2):
@@
{listing}
[h] Overcooked Layout Generator {minted}[fontsize=]python def generate_random_layout(seed, params): rng = random.Random(seed) for attempt in range(params.max_attempts): h = rng.randint(*params.h_range) w = rng.randint(*params.w_range) grid = init_floor_with_border(h, w)
# 1. Interactive tiles for sym in [GOAL, POT, ONION_PILE, PLATE_PILE]: if not place_random(grid, sym, rng.randint(1, 2), rng): break # restart
# 2. Extra walls to hit density n_target = round(params.wall_density * (h-2)*(w-2)) n_add = n_target - count_unpassable(grid) if not place_random(grid, WALL, n_add, rng): continue # restart
# 3. Agents if not place_random(grid, AGENT, params.n_agents, rng): continue
# 4. Validate if evaluate_grid(to_string(grid)): return to_string(grid)
@@ A.3 Layout Validator
Let $G$ be an $h\times w$ character matrix with symbols $\{\texttt{W},\texttt{X},\texttt{O},\texttt{B},\texttt{P},\texttt{A},\textvisiblespace\}$ for walls, delivery, onion pile, plate pile, pot, agent, and floor. Interactive tiles are =$\mathcal{I}=\{\texttt{X},\texttt{O},\texttt{B},\texttt{P}\}$, and unpassable tiles $\mathcal{U}=\mathcal{I}\cup\{\texttt{W}\}.$
@@ 10. R10
Listing A.3 shows a condensed version of the validator.
{listing}
[h] Condensed Layout Validator. {minted}[fontsize=]python def validate(grid_str): g = [list(r) for r in grid_str.splitlines()] h, w = len(g), len(g[0])
# R1–R3 omitted for brevity …
# Depth-first search from a start cell def dfs(i, j, seen): if (i, j) in seen or g[i][j] in UNPASSABLE_TILES - AGENT: return seen.add((i, j)) for di, dj in ((1,0),(-1,0),(0,1),(0,-1)): dfs(i+di, j+dj, seen)
# Agents and family reachability a1, a2 = [(i, j) for i,r in enumerate(g) for j,c in enumerate(r) if c == AGENT] reach1, reach2 = set(), set() dfs(*a1, reach1); dfs(*a2, reach2)
# Helper: reachable(S, reach) def any_reach(symbols, reach): return any(g[i][j] in symbols for i,j in reach)
# R5–R7 if not any_reach(ONION_PILE, reach1|reach2): return False if not any_reach(POT, reach1|reach2): return False if not any_reach(GOAL, reach1|reach2): return False
# R8–R10 (usefulness & hand-off) def useful(reach_me, reach_other): # direct or shared-wall hand-off for i,j in reach_me: if g[i][j] in INTERACTIVE_TILES: return True if g[i][j] == FLOOR and any( (abs(i-i2)+abs(j-j2) == 1 and g[i2][j2] == WALL) for i2,j2 in reach_other): return True return False
if not useful(reach1, reach2): return False if not useful(reach2, reach1): return False return True
@@ B.1 Network Architecture
  • MLP (default) : observation tensor is flattened to a vector and passed through $\{2,3\}$ fully-connected layers of width $128$.
[…]
@@
[…]
@@ B.2 Hyperparameters
Table 4: Fixed hyper-parameters. All experiments use dense reward shaping, two agents, and IPPO unless noted. CL coefficients $\lambda$ refer to the regularisation strength passed to each method.
@@
Optimisation (IPPO)
@@
Importance episodes / steps $5$ / $500$
Regularise critic / heads No / No
@@
Appendix C Extended Results
In this section, we provide additional details, results, and analysis of our experiments.
C.1 Difficulty Levels
Higher difficulty levels pose greater challenges for both learning and retention. As the grid size and obstacle density increase, the environment becomes more complex: interactable items are farther apart, and navigation paths are longer and more convoluted. This increases the number of steps required to complete a recipe, making the overall task harder to memorize and reproduce. Higher-level layouts also add demands for plasticity and transfer. The larger layout space introduces greater variability between tasks, making it harder to reuse learned behavior. These factors collectively lead to lower performance as difficulty increases, as shown in Figure 7.
@@ C.1 Difficulty Levels
C.2 Network Plasticity
C.2.1 Metrics
@@
where 00 indicates perfect retention. We report the mean of Eq. (4) over repetitions and seeds.

### v2

@@ A.2 Procedural Kitchen Generator
Given a random seed and user-selectable parameters (number of agents $n_{a}$, layout height range $h_{\min},h_{\max}]$, layout width range $w_{\min},w_{\max}]$, and wall-density $\rho$), the goal is to emit a solvable grid string $G$ representing the Overcooked environment.
@@ A.2.1 Notation
Let $h,\,w\sim\text{UniformInt}(h_{\min},h_{\max}),\;\text{UniformInt}(w_{\min},w_{\max}),$ and denote by $\mathcal{C}=\{(i,j)\mid 1\leq i\leq h-2,\;1\leq j\leq w-2\}$ the set of internal cells (outer walls excluded). Its cardinality is $N_{\mathrm{int}}=(h-2)(w-2)$. An unpassable cell contains either a hard wall (#) or an interactive tile; we write $N_{\mathrm{unpass}}(G)$ for the number of such cells in $G$.
@@ A.2.2 Algorithm
The generator performs the following loop until a valid grid is produced (Listing 10):
@@
Figure 10: Overcooked Layout Generator
def generate_random_layout(seed, params):
rng = random.Random(seed)
for attempt in range(params.max_attempts):
h = rng.randint(*params.h_range)
w = rng.randint(*params.w_range)
grid = init_floor_with_border(h, w)
[…]
for sym in [GOAL, POT, ONION_PILE, PLATE_PILE]:
if not place_random(grid, sym, rng.randint(1, 2), rng):
break # restart
[…]
n_target = round(params.wall_density * (h-2)*(w-2))
n_add = n_target - count_unpassable(grid)
if not place_random(grid, WALL, n_add, rng):
continue # restart
[…]
if not place_random(grid, AGENT, params.n_agents, rng):
continue
[…]
if evaluate_grid(to_string(grid)):
return to_string(grid)
@@ A.3 Layout Validator
Let $G$ be an $h\times w$ character matrix with symbols $\{\texttt{W},\texttt{X},\texttt{O},\texttt{B},\texttt{P},\texttt{A},\textvisiblespace\}$ for walls, delivery, onion pile, plate pile, pot, agent, and floor. Interactive tiles are $\mathcal{I}=\{\texttt{X},\texttt{O},\texttt{B},\texttt{P}\}$, and unpassable tiles $\mathcal{U}=\mathcal{I}\cup\{\texttt{W}\}.$
@@ 10. R10
Listing 11 shows a condensed version of the validator.
Figure 11: Condensed Layout Validator.
def validate(grid_str):
g = [list(r) for r in grid_str.splitlines()]
h, w = len(g), len(g[0])
[…]
def dfs(i, j, seen):
if (i, j) in seen or g[i][j] in UNPASSABLE_TILES - {AGENT}:
return
seen.add((i, j))
for di, dj in ((1,0),(-1,0),(0,1),(0,-1)):
dfs(i+di, j+dj, seen)
[…]
a1, a2 = [(i, j) for i,r in enumerate(g)
for j,c in enumerate(r) if c == AGENT]
reach1, reach2 = set(), set()
dfs(*a1, reach1); dfs(*a2, reach2)
[…]
def any_reach(symbols, reach):
return any(g[i][j] in symbols for i,j in reach)
[…]
if not any_reach({ONION_PILE}, reach1|reach2): return False
if not any_reach({POT}, reach1|reach2): return False
if not any_reach({GOAL}, reach1|reach2): return False
[…]
def useful(reach_me, reach_other):
# direct or shared-wall hand-off
for i,j in reach_me:
if g[i][j] in INTERACTIVE_TILES: return True
if g[i][j] == FLOOR and any(
(abs(i-i2)+abs(j-j2) == 1 and g[i2][j2] == WALL)
for i2,j2 in reach_other):
return True
return False
[…]
if not useful(reach2, reach1): return False
return True
@@ B.1 Network Architecture
  • MLP (default) : observation tensor is flattened to a vector and passed through 2 fully-connected layers of width $128$.
  • CNN : three $32$-channel convolutions with kernel sizes 5×5, 3×3, ${\times}5,\,3{\times}3$${\times}3$ feed a $64$-unit projection, followed by a single $128$-unit dense layer.
@@
  • Multi-head outputs : if use_multihead is set, each head holds a distinct slice of logits/values for every task ($\texttt{num\_tasks}=|\mathcal{T}|$). The correct slice is selected with the cheap tensor reshape in choose_head.
@@ B.2 Hyperparameters
Table 6: Fixed hyper-parameters. All experiments use dense reward shaping, two agents, and IPPO unless noted. CL coefficients $\lambda$ refer to the regularization strength passed to each method.
@@
Optimization (IPPO)
Activation ReLU
Optimizer Adam (Optax)
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999)(0.9,\ 0.999)
Adam $\epsilon$ $10^{-5}$
Weight decay none
@@
Importance episodes / steps 55 / $500$
Regularize critic / heads No / No
AGEM Memory size 100 000 transitions
AGEM Sample size (per proj.) 1024
@@
Random seeds $\{1\,..\,5\}$
Appendix C EWC vs. Online EWC
Figure 12: Comparison of EWC and Online EWC across all difficulty levels in terms of forward transfer and forgetting. Each point denotes a method’s performance at a given level. Online EWC consistently exhibits higher plasticity (less-negative or positive, particularly at Level 3, while EWC achieves notably lower forgetting on all levels.
EWC accumulates importance over all past tasks and penalizes drift along high-Fisher directions with a fixed quadratic. Online EWC maintains a running, exponentially decayed Fisher, emphasizing recent tasks and relaxing old constraints. Both use the same heads, meaning that the penalty acts on the shared trunk. When layouts are small, not only are the tasks easier to learn, but the same features are more likely to work across tasks. Strong anchoring preserves those features, curbing forgetting and yielding a higher average score. The stability–plasticity trade-off is favorable because plasticity demands are modest. This trade-off is visualized in Figure 12, where Online EWC demonstrates higher plasticity at the cost of increased forgetting, while EWC excels in stability but struggles to adapt on Level 3. Level 3 forces longer paths, bottlenecks, and role specialization, which require larger representational updates. EWC’s cumulative constraints over-tighten the trunk and slow adaptation, while Online EWC’s decay frees capacity for those shifts, so it learns the hard tasks more effectively. The multiple output heads alone are not enough. They isolate outputs, but the penalty sits on the shared backbone. When the trunk needs to be to rewired for new Level 3 tasks, EWC resists too much, while Online EWC allows it more. Moreover, credit assignment is noisier on Level 3 due to sparser effective signals and longer horizons. A single, stale Fisher snapshot can misdirect EWC’s penalty. The rolling estimate in Online EWC smooths that noise and tracks the current regime more closely.
Appendix D Partially Observable MEALs
To more closely mimic the constraints faced by real-world agents, we introduce a direction-aware egocentric observation setting. Each agent perceives a rectangular window centered on itself, with tiles outside this window masked. The window is anisotropic with respect to the agent’s heading: we separate forward, side, and rear extents, which increase with difficulty (Table 7). This scaling is intentionally balanced with the overall environment design: as the grid size grows with difficulty, the perceptual window also expands to maintain a comparable challenge-to-information ratio. Consequently, the tasks become POMDPs, where exploration, memory (e.g., recurrent state), and implicit/explicit coordination provide tangible benefits. In particular, Level 1 removes rear context entirely, Level 2 extends the look-ahead by one tile, and Level 3 adds both longer look-ahead and rear visibility, reducing blind spots while preserving partial observability (Figure 13).
Table 7: Field-of-view specification for the partially observable MEAL variant. Window size and directional extents scale with difficulty.
@@ Appendix D Partially Observable MEALs
[…]
Medium 8–9 2 1 0 $\times3$
[…]
(a) Level 1: $2{\times}3$ window.
(b) Level 2: $3{\times}3$ window.
(c) Level 3: $3{\times}5$ window.
Figure 13: Egocentric observation windows by difficulty. Visibility grows with difficulty but remains partial, preserving the need for exploration and memory.
Appendix E Curriculum Learning
In all training settings, agents consistently struggle on Level 3 tasks with large grids. Curriculum learning has been shown to improve final performance on difficult tasks by gradually increasing task complexity [Bengio et al., 2009, Narvekar et al., 2020, Portelas et al., 2020]. We investigate whether a simple difficulty-based curriculum can help agents better learn harder MEAL tasks under the same data budget. To this end, we design a curriculum sequence where each difficulty level contributes an equal number of tasks. Specifically, we sample 5 layouts each from Level 1 (easy), Level 2 (medium), and Level 3 (hard), and present them in ascending order of difficulty (layouts 1–5, then 6–10, then 11–15). As a baseline, we compare with a default sequence that trains on 15 hard (Level 3) layouts without any prior exposure to easier tasks. Performance is evaluated based on the normalized average score over the 5 tasks in the sequence of the respective difficulty.
Table 8: Curriculum vs. default training under an equal data budget. We report the average score over the task windows of the respective difficulty.
Strategy Medium (6–10) Hard (11–15)
Default 0.693 $\pm$ 0.147 0.328 $\pm$ 0.238
Curriculum 0.668 $\pm$ 0.152 0.653 $\pm$ 0.181
The results in Table 8 show no statistically significant difference between the two strategies on Level 2, given the high variance. However, on Level 3, the curriculum strategy nearly doubles performance. A plausible explanation is that, under curriculum training, the agent first experiences 5 easy and 5 medium tasks, where it receives denser reward signals and more frequent successes. This exposure likely builds useful priors and stabilizes learning, improving adaptation to harder tasks later. In contrast, the default strategy trains only on hard tasks throughout the sequence, where exploration is more challenging and initial rewards are more difficult to obtain, leading to weaker performance overall.
Appendix F Difficulty Levels
Higher difficulty levels pose greater challenges for both learning and retention. As the grid size and obstacle density increase, the environment becomes more complex: interactable items are farther apart, and navigation paths are longer and more convoluted. This increases the number of steps required to complete a recipe, making the overall task harder to memorize and reproduce. Higher-level layouts also add demands for plasticity and transfer. The larger layout space introduces greater variability between tasks, making it harder to reuse learned behavior. These factors collectively lead to lower performance as difficulty increases, as shown in Figure 14. EWC performs reliably on Level 1, successfully learning new tasks while retaining performance on learned ones. On Level 2, learning remains effective up to around the midpoint of the sequence, after which progress slows and variance across runs increases markedly. In Level 3, EWC struggles from the outset, as performance is poor even on early tasks, and learning new tasks stalls entirely after the 15th task in the sequence.
@@ Appendix F Difficulty Levels
Appendix G Network Plasticity
G.1 Metrics
@@
where 0 indicates perfect retention. We report the mean of Eq. (5) over repetitions and seeds.
@@ G.2 Training Curves
Appendix H Designated Roles
In Overcooked, agents are identical in their capabilities and attributes. However, in many real-world scenarios, autonomous agents either 1) possess different physical properties or 2) are functionally identical but are expected to fulfill distinct, complementary roles to cooperate effectively for a common goal. To capture this dimension in MEAL, we design a heterogeneous agent setting with designated roles.
In this variant, two agents are randomly assigned one of the two predefined roles at the start of each task: chef and waiter. The chef is responsible for preparing the soup by loading onions into the pot, but cannot pick up plates. The waiter handles dish delivery but cannot pick up onions. This enforces complementary capabilities, meaning neither agent can complete the full recipe alone, meaning that successful catering requires coordinated role execution and adaptation. Note that the roles are sampled per task and may switch across tasks, making continual learning essential.
Table 9: Homogeneous vs. heterogeneous (designated roles) 2-agent training results over Level 1 20-task generated sequences using shared rewards and IPPO in combination with EWC.
Setting $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
Homogeneous 0.90 $\pm$ 0.04 0.01 $\pm$ 0.01 0.20 $\pm$ 0.08
Heterogeneous 0.68 $\pm$ 0.09 0.03 $\pm$ 0.02 0.05-0.05 $\pm$ 0.09
We evaluate this setting over 20-task Level 1 sequences using EWC with IPPO under shared rewards. Table 9 compares the heterogeneous setup to the default homogeneous setting. We observe a clear performance drop in the role-restricted setting, as throughput decreases when agents are limited to certain actions and cannot flexibly switch between tasks. Another factor is asymmetric step costs: in many layouts, loading the pot with 3 onions takes more steps than a single plate-and-deliver trip, making the chef the throughput bottleneck. Generalization also suffers as agents struggle to transfer knowledge when their roles change across tasks, since skills learned in one role do not apply to the other. This role-switching dynamic further exacerbates forward transfer challenges in continual learning.
Appendix I Common Pitfalls
Despite shared rewards and simple layouts, learned policies frequently fall into recurring failure modes that throttle throughput and coordination. Figure 16 illustrates three such patterns we observe consistently across layouts and levels.
(a) Single-pot fixation. All agents are clustered around a single pot, waiting for it to finish cooking, while ignoring a ready soup in the bottom pot.
(b) Deadlock. The red agent tries to place an onion into the pot, but is blocked by the blue agent, who cannot move aside.
(c) Role collapse. One agent completes the pipeline solo while the other wanders or idles. The policy settles on a local minimum.
Figure 16: Qualitative failure modes observed in Overcooked. All behaviors stem from inadequate coordination, limited exploration, or insufficient role allocation.
Appendix J Extended Results
In this section, we provide additional experimental results. Tables 10, 11, 12 add 95% confidence intervals to the main baseline results. Figure 19 depicts the per-task evaluation curves of Level 1. Figure 20 illustrates forward transfer.
Figure 17: Average Normalized Score curves on Level 2.
Figure 18: Average Normalized Score curves on Level 3.
Table 10: Level 1 baseline results with confidence intervals.
Method $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
[…]
Table 11: Level 2 baseline results with confidence intervals.
Method $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
[…]
Table 12: Level 3 baseline results with confidence intervals.
Method $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
[…]
Figure 19: The evaluation curves of Level 1 illustrate the extent of forgetting across tasks. FT suffers from clear catastrophic forgetting: once the agent transitions to a new task, performance on the previous task collapses immediately. EWC and L2 display near-perfect retention.
Figure 20: Forward transfer on Level 1. The green shaded areas depict positive transfer compared to the IPPO baseline, and the red shaded areas show negative transfer.